OpenStudy (anonymous):

PLEASEEE HELP PLEASSE URGENT.(MEDAL+FAN) 1.What is the common difference of a 57-term arithmetic sequence where the first term is -25 and the sum is 17,727? (1 points) 11 11.5 12 12.5 2.What is the sum of a 14-term arithmetic sequence where the last term is 30 and the common difference is -5? (1 points) 875 812.5 455 422.5 3.When finding the sum of the arithmetic sequence 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Use complete sentences to describe how the -3 was calculated and what it means.

OpenStudy (anonymous):

@satellite73 @hartnn @Hero , Can Yall Please Help & Explain Please This IS Urgent

hartnn (hartnn):

which one ?

OpenStudy (anonymous):

1

hartnn (hartnn):

\(\Large S_n = (n/2) [2a_1 +(n-1)d]\) here, Sn = 17727 a1 = -25 n = 57 just plug in values! only d is unknown

OpenStudy (anonymous):

so it'sgoing to b 17727=(57/2)[2(-24)+(57-1)d]

hartnn (hartnn):

a1 is actually -25, not -24

OpenStudy (anonymous):

oh yea so i solve it now rite

hartnn (hartnn):

yes! simplify and get the value of d :)

OpenStudy (anonymous):

I Can't put it into thecalculator norsolve it like its set up

hartnn (hartnn):

multiply both sides by 2/57 first

hartnn (hartnn):

left side = 17727*2/57 = ...

OpenStudy (anonymous):

622

hartnn (hartnn):

correct right side = -50 +56 d now add 50 on both sides

hartnn (hartnn):

622 +50 = 56d can you find d now ?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

good :)

OpenStudy (anonymous):

d=12

hartnn (hartnn):

absolutely correct :)

OpenStudy (anonymous):

Thank You So Much. But Can U help me set up the equations for the 2.PLEASE ^.^

hartnn (hartnn):

sure

hartnn (hartnn):

we first need to find the 1st term a1 use the n'th term formula \(\Large a_n = a_1 + (n-1)d\) here, n= 14 an =30 d =-5 plug in! find a1, its the only unknown

OpenStudy (anonymous):

30=a1+(14-1)-5

hartnn (hartnn):

yup don't miss the brackets 30=a1+(14-1)(-5)

hartnn (hartnn):

30 = a1 - 13*5 can u find a1 from here ?

OpenStudy (anonymous):

Yea. I Think let me try

OpenStudy (anonymous):

I Got 95 :/

hartnn (hartnn):

a1 = 95 is correct :)

hartnn (hartnn):

now use the same sum formula i gave you before a1 = 95 d=-5 n=15 plug these in and find the sum!

hartnn (hartnn):

try this all the way by yourself first :)

OpenStudy (anonymous):

k

hartnn (hartnn):

15/2 (2*95 + 13*(-5)) did you get this after plugging in ?

hartnn (hartnn):

sorry 14/2 (2*95 + 13*(-5))

hartnn (hartnn):

ohh...i wrote n=15,its actually n=14 because we need sum of 14 terms

OpenStudy (anonymous):

K Thank YOU SOO SOOO MUCH. YOU THE BEST ! I GOT 875. THANKS A TON AGAIN

hartnn (hartnn):

875 is correct :) welcome ^_^